The coefficient of x4 in is :







 


 


 

Grade:12th Pass

2 Answers

aakriti goel
13 years ago

d ans is d 1st option

Nehal Wani
13 years ago

 

 

=10r(x/2)10-r(-3/x2)r

=10r(x/2)10-r(-3)r(x-2)r

10-r-2r=4

10-3r=4

r=2

Coefficient Of x4 in

 

 

=102(1/2)10-2(-3)2

=45*9/28

=405/256